In a large population of adults, the mean IQ and standard deviation are

\mu =112;\sigma =20

respectively.

Suppose 200 adults are randomly selected for a market research campaign.  What is the probability that the sample mean is less than 110?
Answer #1

Solution :

Given that ,

mean = \mu = 112

standard deviation = \sigma = 20

n = 200

\mu\bar x = 112

\sigma\bar x =\sigma  / \sqrt n = 20/ \sqrt 200=1.4142

P(\bar x <110 ) = P[(\bar x - \mu \bar x ) / \sigma \bar x < (110 -112) / 1.4142]

= P(z < -1.41)

Using z table  

= 0.0793

probability=0.0793